Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 8,400 units, 4/5 completed 22,512
31 Direct materials, 151,200 units 302,400 324,912
31 Direct labor 85,880 410,792
31 Factory overhead 48,310 459,102
31 Goods finished, 153,300 units 444,234 14,868
31 Bal. ? units, 2/5 completed 14,868

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit. $
2. Conversion cost per equivalent unit. $
3. Cost of the beginning work in process completed during March. $
4. Cost of units started and completed during March. $
5. Cost of the ending work in process. $
Answer #1

Answer a(1):

Direct materials cost per equivalent unit = cost of direct material added / quantity of direct material added

= 302,400 / 151,200

= $2

Direct materials cost per equivalent unit = $2

Answer a(2):

Finished goods cost per equivalent unit = 444,234 / 153,300 = $2.90

Conversion cost per equivalent unit = Finished goods cost per equivalent unit - Direct materials cost per equivalent unit.

= $2.90 - $2

= $0.90

Conversion cost per equivalent unit = $0.90

Answer a(3):

Beginning inventory had 8,400 units and 4/5 of conversion was completed.

During March balance 1/5th conversion of 8400 units was completed. Cost for the same = 1/5 * 8400 * 0.90 = $1512

Cost of the beginning work in process completed during March = 22512 + 1512 = $24,024

Answer a(4):

Cost of units started and completed during March = cost of finished goods transferred - Cost of the beginning work in process completed during March

= 444,234 - 24024

= $420,210

Cost of units started and completed during March = $420,210

Answer a(4):

Cost of the ending work in process = $14,868